OpenStudy (anonymous):

What is the slope of the tangent line to the graph of f(x) = 5x^2 - 4x at P(a, f(a))?

jimthompson5910 (jim_thompson5910):

f(x) = 5x^2 - 4x f'(x) = 2*5x^1 - 1*4x^0 f'(x) = 10x - 4 ------------------ Slope of tangent line at P(a, f(a)) is f'(a) = 10a-4